OpenStudy (anonymous):

MEDALS GIVEN Without dividing, do you expect seven-eighths divided by eleven-sixteenths to be greater than or less than 1? Explain.

OpenStudy (anonymous):

7*16=112 and 8*11 = 88, therefore I would say yes it would be greater than 1.

OpenStudy (anonymous):

suppose \(x> b\) and neither are 0. then \(\left|\frac{x}{b}\right|>1\) and \(\left|\frac{b}{x}\right|<1\)

OpenStudy (anonymous):

I could also say that you are taking a number slightly below 1 and dividing it by another number even smaller than the one originally had, and that would give you a number greater than 1. Like if I started with three quarters and I divide it by 1/2 my number is larger than one, but that does not work for 1/2 divided by 3/4. Because my original was greater than the number i divided it by.